# arXiv:2603.01359v1[eess.SY] 02 Mar 2026

### v1

@@ 2.3 NExT-LF
Summarising, NExT can retrieve IRFs from output-only data. IRFs are the time domain counterpart of the FRFs, such that they can be brought into the frequency domain by means of an FTT and then become a suitable input for frequency domain methods, such as the LF. In this work, NExT-LF follows the workflow shown in Figure˜1.
@@ 3 The eXperimental BeaRDS 2 Flexible Wing Spar Model
The experimental dataset employed in this work is sourced from Dessena et al. (2022) [^1] The experimental and benchmark data can be retrieved from https://doi.org/10.17862/cranfield.rd.19077023, which also reports benchmark EMA results. The spar is excited for 20 m using a bandwidth-limited random input in the range 2-400 Hz, with an RMS amplitude of 0.305 g. The excitation is applied using a Data Physics Signal Force™ modal shaker operated under closed-loop control via DP760™ software. As shown in Figure˜2, the accelerometers are arranged in a 4 by 2 grid to measure vertical acceleration, thereby capturing both flapwise (out-of-plane bending in the $x$-$z$ plane) and pitching (torsion around the $x$ axis) responses. The signals were sampled at $f_{s}=$ 5120 Hz using a National Instruments cDAQ-9178 and saved via an in-house LabVIEW programme [^2] At the time the first author was affiliated with Cranfield University.. The recorded acceleration time series are then converted to ms-2, band-pass filtered between 3.25 and 85 Hz to suppress low-frequency drift, and band-stop filtered between 49.5 and 50.5 Hz to remove mains-related interference [^3] In the UK, the mains electricity AC frequency is 50 Hz.. The upper passband limit is set to 85 Hz, since all modes of interest lie below this frequency. Further details on the experimental setup and instrumentation used in the test are avaible in Dessena et al. (2022) and additional details on a nonlinear vibration testing campaign on XB-2 and components can be found in Dessena et al. (2025b).
Using the eight output-channel time histories, the response spectra can be estimated in terms of power spectral density (PSD). This is computed in MATLAB via the built-in pwelch function [^4] https://uk.mathworks.com/help/signal/ref/pwelch.html. Prior the PSD estimation, the signal is downsampled to 256 Hz. Then, the resulting PSD estimates, as shown in Figure˜3, are obtained by adopting 2048 discrete Fourier transform points, an overlap of 256 samples, and a Hamming window of length 1024 to reduce variance and provide a smoother spectral estimate.
@@
After this step, the identification procedure begins by providing NExT with the eight measured output acceleration time series, from which IRFs discretised over 2048 points are obtained. In this work, the acceleration signal from accelerometer #8 (see Figure˜2) is used as the sole reference channel for the NExT algorithm. This choice is motivated by the observation that the corresponding NExT-derived FRFs, as shown in Figure˜4), present the three expected resonance peaks more clearly (higher amplitude) than when other reference channels are used.
@@
To limit the influence of outliers and non-physical modes, the identification is repeated across multiple model orders $k$, with $k\in[6,50]$, and stabilisation diagrams are used to retain only stable modes. The range of $k$ was chosen to obtain a comparable count of stable modes across the methods. As discussed later, this condition could not be met for ERA. The stability assessment considers $\omega_{n}\in[0,83]\,\mathrm{Hz}$ with a stability parameter of $0.5\%$, and $\zeta_{n}\in[0.01,0.03]$ with a stability parameter of $5\%$. A mode is accepted as stable only when the criteria are satisfied for five consecutive identifications. $\bm{\phi}_{n}$ consistency is enforced by requiring a Modal Assurance Criterion (MAC) value not smaller than $0.95$ among the subsequent ($k$-wise) identified modes. The stabilisation diagrams for NExT-LF, NExT-ERA, and SSI-CVA are reported in Figure˜4.
@@
As shown in Figure˜5, the NExT-LF and SSI-CVA results exhibit markedly stronger stabilisation than the NExT-ERA case, with the difference being most evident in the damping estimates $\zeta_{n}$. In particular, NExT-ERA yields virtually no modes that stabilise in damping. This behaviour is consistent with the identified $\omega_{n}$ and $\zeta_{n}$ values reported in Tables˜1 and 2, respectively.
@@
The $\omega_{n}$ identified using NExT-LF, NExT-ERA, and SSI-CVA are in close agreement with the benchmark results reported in Dessena et al. (2022), with the largest discrepancy being $3.42\%$ for $\omega_{1}$ obtained via NExT-ERA. In contrast, the $\zeta_{n}$ in Table˜2 show noticeably larger differences. Specifically, the absolute error in $\zeta_{1}$ exceeds $50\%$ for both NExT-LF and NExT-ERA, with NExT-LF providing a smaller deviation than NExT-ERA, whereas SSI-CVA limits the absolute error to approximately $17\%$. For $\zeta_{2}$ and $\zeta_{3}$, the NExT-LF estimates deviate by about $7.3\%$ in absolute terms, while NExT-ERA produces larger errors, namely $11.4\%$ and $7.46\%$, respectively. Overall, SSI-CVA yields damping estimates that are most consistent with the reference identification, although the absolute error for $\zeta_{1}$ remains above $17\%$.
After discussing the identifications of $\omega_{n}$ and $\zeta_{n}$, the results for the $\bm{\phi}_{n}$ are now considered. Consistently with the frequency estimates, the mode shapes identified from the output-only measurements show excellent agreement with the benchmark, with the diagonal MAC values being equal, or very close, to unity. Figure˜6 presents the $\bm{\phi}_{n}$ obtained via NExT-LF, NExT-ERA, and SSI from output-only data, overlaid with the reference $\bm{\phi}_{n}$ reported in Dessena et al. (2022). The identified $\bm{\phi}_{n}$ closely reproduce the benchmark results.
Figure 6: $\bm{\phi}_{n}$ of the XB-2 wing spae identified using the output-only methods, overlaid with the benchmark results (retrieved from Dessena et al. (2025a)).
@@ 4 Airbus Helicopters H135 Bearingless Main Rotor Blade
In order to generalise the findings of the previous sections, a second benchmark aeronautical structure should be considered. For this aim an Airbus Helicopters H135 BMR blade is chosen. The Airbus Helicopters H135, shown in Figure˜7(a), is a lightweight rotorcraft capable of a maximum forward flight speed of 136 kts, a maximum range of 342 nm, and a maximum take-off mass of 3100 kg. In this work, the focus is on its BMR blade, which is made of the pitch control cuff with an elastomeric lead-lag damper, an aerofoil section and a flexbeam. This arrangement allows it to incorporate the conventional blade mechanical flapping, lagging, and torsional hinges into a single structural arrangement. More information on the Airbus Helicopters H135 BMR blade is available in Bansemir and Müller (1997). From a geometric perspective, the BMR blade has an equivalent chord of 0.288 m and spans 4.99 m from the hub attachment. The H135 BMR blade specimen considered here, see Figure˜7(b) is an unserviceable [^5] In the aircraft maintenance sense. item, still in relatively good shape, which has been dynamically tested, via AVT, in the former Cranfield University project BladeSense Weber et al. (2021).
@@
Figure 7: (a) Airbus Helicopters H135 helicopter in Western Counties Air Operations Unit livery (UK – retrieved from Felce (2011)) and (b) the test setup of the ambient vibration testing on the Airbus Helicopter H135 BMR blade (adapted from Weber et al. (2021)).
The AVT was carried out in the aeroelastic laboratory of the Sir Peter Gregson Aerospace Integration Research Centre at Cranfield University, under controlled temperature and humidity conditions. The stationary, i.e. not revolving, blade is clamped at its end and instrumented with 9 nine triaxial accelerometers (A1-9), recording in-plane ($z$-axis) and out-of-plane transverse ($y$-axis) acceleration response from ambient vibration666xx-axis accelerations are recorded from A9, but are disregarded in this study.. The accelerometer types are shown in Figure˜7(b) and their layout and spanwise position in Figure˜8. The accelerometers spanwise locations in Figure˜8 are determined graphically from the Figures in Weber et al. (2021). Additionally, note that   highlights the 4 PCB Model 356B18 accelerometers used, while   and   identify, respectively, the 3 PCB Model 356A16 and 2 PCB Model 356A45 accelerometers.
@@
From the 9 accelerometers, 18 acceleration time series are recorded at a sampling frequency $f_{s}=$ 2560 Hz for 600 s, using a similar setup to what is described for the XB-2 spar in Section˜3, involving a National Instruments cDAQ-9178 and then in-house LabVIEW programme. Prior to identification, the signals are decimated to 256 Hz, demeaned, detrended, and bandpass-filtered between 0.4 and 100 Hz to exclude potential low- and high-frequency drifts with the bandpass [^7] https://uk.mathworks.com/help/signal/ref/bandpass.html built-in MATLAB function. From here, a PSD of the 18 channels and the ANPSD are found to qualitatively assess the quality of the response . The PSDs and ANPSD, shown in Figure˜9 are calculated using Welch’s method over 2048 discrete points with a Hamming window of length 2048 with an overlap of 256.
Figure 9: PSDs and ANPSD from the signals recorded from accelerometers A1-9 on the irbus Helicopter H135 BMR blade.
As it is clear from Figure˜9, this case poses a tougher identification challenge than the XB-2 spar, as the AVT amplitudes are much smaller than those produced by shaker excitation, resulting in lower modal amplitudes. As for the other cases, the identification procedure starts by obtaining the IRF from NExT, discretised over 4098 points, using two reference channels, A1 in $x$ and A4 in $x$. This is necessary in this case, as using a reference channel for each direction increases the likelihood of obtaining accurate modal parameter estimates Caicedo (2011). The obtained IRFs are fed directly into ERA, while they are transformed to the frequency domain via an FFT discretised over 2048 points for use with the LF. The SSI-based identification is carried out directly from the time-domain data.
The identification is carried at a given order range $k\in[22,130]$ and stabilisation diagrams are used to exclude spurious and non physical modes. For this aim, stabilisation parameters, remaining the same across NExT-LF, NExT-ERA, and SSI-CVA, are defined. The stabilisation assessment is carried out over $f_{n}\in[0,100]\,$ Hz and $\zeta_{n}\in[0.003,0.03]$. Frequency stability is enforced with a relative tolerance of $0.5\%$ (i.e., $\Delta\omega_{n}/\omega_{n}\leq 0.005$), while damping stability is checked with a $5\%$ tolerance (i.e., $\Delta\zeta/\zeta\leq 0.05$). An additional absolute criterion on the frequency variation is also applied, $|\Delta\omega_{n}|\leq\epsilon$, with $\epsilon=0.1$. A mode is accepted as stable only if these criteria are satisfied for three consecutive identifications. The consistency of the $\bm{\phi}_{n}$ is enforced by requiring a MAC not smaller than $0.95$ among the subsequent stable identified modes. The stabilisation diagrams obtained from the NExT-LF, NExT-ERA, and SSI-CVA identification reported above are shown in Figure˜10.
@@
The extracted modal parameters, $\omega_{n}$ and $\zeta_{n}$, are presentend in Table˜3 with results from the literature. These include results from AOMA Mugnaini et al. (2022), containing also $\bm{\phi}_{n}$, and Weber et al. (2021), including $\omega_{n}$ from a calibrated FEM and $\omega_{n}$ and $\zeta_{n}$ from a separate experimental configuration, based on a series of GVTs conducted using a random-on-random excitation waveform. Please note that in Table˜3, flapping refers to out-of-plane bending (within the xxyy plane), lagging to in-plane bending (within the xxyy plane), pitching to torsional rotations about the $x$-axis, and coupled to a mode exhibiting a combination of two or more of the dominant motions described above.
@@
From Table˜3, it is important to note that NExT-LF identifies two additional modes within the frequency range of interest, compared with the previous identification reported for AOMA Mugnaini et al. (2022) and the experimental results in Weber et al. (2021). These modes are located at $19.69~\mathrm{Hz}$ and $59.14~\mathrm{Hz}$, respectively. In addition, SSI-CVA identifies a further mode at $68.09~\mathrm{Hz}$. As shown later, these modes are all coupled in the $z$- and $y$-axis displacements, with the suspicion that the 2nd coupled mode may also include a torsional component; however, this cannot be fully confirmed due to the adopted instrumentation layout. Nevertheless, the coupled character of these modes provides a plausible explanation for why they are not identified by the simplified FEM reported in Weber et al. (2021). Furthermore, it is worth noting that NExT-LF finds more modes than NExT-ERA (13 vs 9), and the same number as SSI-CVA.
@@
In terms of the identified natural frequencies $\omega_{n}$, the agreement across the different output-only identification methods is generally very good. With respect to the reference values obtained via AOMA Mugnaini et al. (2022), the relative differences remain below approximately $3\%$ for all coincident modes, the largest deviation being observed for NExT-ERA at the 10th mode. A similar level of agreement is observed when comparing with the experimental results reported in Weber et al. (2021). The consistency of the identified $\omega_{n}$ across NExT-LF, NExT-ERA, and SSI-CVA confirms the robustness of the proposed framework in terms of frequency estimation, even for higher-order flapping and lagging modes.
@@
However, the same level of consistency is not observed for the damping ratios $\zeta_{n}$. In contrast to $\omega_{n}$, the identified $\zeta_{n}$ values exhibit a significantly larger dispersion across methods. In particular, relative differences with respect to AOMA Mugnaini et al. (2022) frequently exceed several tens of percent and, in some cases, are considerably larger, especially for the lower-frequency modes. This behaviour is expected, as damping estimation in operational modal analysis is well known to be more sensitive to noise, limited measurement duration, and modelling assumptions than frequency estimation. Moreover, small absolute differences in $\zeta_{n}$ can translate into large relative deviations when the reference damping level is low.
@@
Overall, the results indicate that while the identification of $\omega_{n}$ is highly reliable and consistent across the considered approaches, the estimation of $\zeta_{n}$ remains more method-dependent. However, to fully assess the performance of NExT-LF, the identified $\bm{\phi}_{n}$ must also be examined. To this end, Table˜4 reports the diagonal MAC values between the $\bm{\phi}_{n}$ identified via NExT-LF and those obtained in Mugnaini et al. (2022).
@@
The MAC results in Table˜4 further contextualise the $\omega_{n}$ and $\zeta_{n}$ comparisons. In particular, NExT-LF achieves very high correlation for several dominant modes (e.g., Modes 1, 3, 4, 8, and 9, with MAC values close to unity), supporting the strong agreement observed for $\omega_{n}$ and indicating that the associated $\bm{\phi}_{n}$ are consistently retrieved. At the same time, the MAC values highlight that the increased modal content identified by NExT-LF comes with a non-uniform level of shape fidelity across all modes: some modes exhibit reduced correlation with respect to Mugnaini et al. (2022), notably Mode 5 (MAC =0.77=0.77) and, more markedly, Mode 7 (MAC =0.44=0.44), as well as very low agreement for Mode 10 (MAC =0.16=0.16) and Mode 14 (MAC =0.09=0.09). This suggests that, although NExT-LF is effective in extracting a larger set of modes than NExT-ERA, a subset of the identified shapes may be more sensitive to the limited instrumentation and to the specific characteristics of the identification procedure, leading to weaker correspondence with the AOMA reference in Mugnaini et al. (2022). Conversely, SSI-CVA shows consistently high MAC values for the modes that can be compared (e.g., Mode 7 and Mode 10), indicating stronger agreement in those cases, albeit without the same increase in the number of extracted modes relative to NExT-ERA. For completeness, the 14 identified $\bm{\phi}_{n}$ are shown in Figure˜11. Note that in the cases where the MAC is more than 0.95, the NExT-LF-identified $\bm{\phi}_{n}$ are shown; otherwise, $\bm{\phi}_{n}$ from Mugnaini et al. (2022) are preferred, with an exception made for mode 12 (3rd coupled), which is retrieved from the SSI-CVA identification.
@@
All $\bm{\phi}_{n}$ shown in Figure˜11 are consistent with the dominant motions reported in Table˜3. Nevertheless, the trajectories associated with the two pitching modes (Figures˜11(g) and 11(n)) cannot be directly inferred from the corresponding plots, as rotational degrees of freedom are not measured on the blade. These modes are therefore classified as pitching based on the FEM evidence reported in Weber et al. (2021). In addition, the three modes newly identified in the present campaign—namely, the coupled modes (Figures˜11(f), 11(k) and 11(l))—do not exhibit a clear dominant displacement direction. For this reason, they are labelled as coupled, at least in terms of combined lagging and flapping components.

### v2

@@ 2.3 NExT-LF
In summary, NExT can retrieve IRFs from output-only data. IRFs are the time domain counterpart of the FRFs, such that they can be brought into the frequency domain by means of an FTT and then become a suitable input for frequency domain methods, such as the LF. In this work, NExT-LF follows the workflow shown in Figure˜1.
@@ 3 The eXperimental BeaRDS 2 Flexible Wing Spar Model
The experimental dataset employed in this work is sourced from Dessena et al. (2022) [^1] The experimental and benchmark data can be retrieved from https://doi.org/10.17862/cranfield.rd.19077023, which also reports benchmark EMA results. The spar is excited for 20 m using a bandwidth-limited random input in the range 2-400 Hz, with an RMS amplitude of 0.305 g. The excitation is applied using a Data Physics Signal Force™ modal shaker operated under closed-loop control via DP760™ software. As shown in Figure˜2, the accelerometers are arranged in a 4 by 2 grid to measure vertical acceleration, thus capturing both flapwise (out-of-plane bending in the $x$-$z$ plane) and pitching (torsion around the $x$ axis) responses. The signals were sampled at $f_{s}=$ 5120 Hz using a National Instruments cDAQ-9178 and saved via an in-house LabVIEW programme [^2] At the time the first author was affiliated with Cranfield University.. The recorded acceleration time series are then converted to ms-2, band-pass filtered between 3.25 and 85 Hz to suppress low-frequency drift, and band-stop filtered between 49.5 and 50.5 Hz to remove mains-related interference [^3] In the UK, the mains electricity AC frequency is 50 Hz.. The upper passband limit is set to 85 Hz, since all modes of interest lie below this frequency. More details on the experimental setup and instrumentation used in the test are available in Dessena et al. (2022); additional details on a nonlinear vibration testing campaign on XB-2 and components can be found in Dessena et al. (2025b).
Using the eight output-channel time histories, the response spectra can be estimated in terms of power spectral density (PSD). This is computed in MATLAB via the built-in pwelch function [^4] https://uk.mathworks.com/help/signal/ref/pwelch.html. Prior to the PSD estimation, the signal is downsampled to 256 Hz. Then, the resulting PSD estimates, as shown in Figure˜3, are obtained by adopting 2048 discrete Fourier transform points, an overlap of 256 samples, and a Hamming window of length 1024 to reduce variance and provide a smoother spectral estimate.
@@
After this step, the identification procedure begins by providing NExT with the eight measured output acceleration time series, from which IRFs are discretised over 2048 points. In this work, the acceleration signal #8 (see Figure˜2) is used as the only reference channel for the NExT algorithm. This choice is motivated by the observation that the corresponding NExT-derived FRFs, as shown in Figure˜4), present the three expected resonance peaks more clearly (higher amplitude) than when other reference channels are used.
@@
To limit the influence of outliers and non-physical modes, the identification is repeated across multiple model orders $k$, with $k\in[6,50]$, and stabilisation diagrams are used to retain only stable modes. The range of $k$ was chosen to obtain a comparable count of stable modes between the methods. As discussed later, this condition could not be met for ERA. The stability assessment considers $\omega_{n}\in[0,83]\,\mathrm{Hz}$ with a stability parameter of $0.5\%$, and $\zeta_{n}\in[0.01,0.03]$ with a stability parameter of $5\%$. A mode is accepted as stable only when the criteria are satisfied for five consecutive identifications. The consistency of $\bm{\phi}_{n}$ is enforced by requiring a Modal Assurance Criterion (MAC) value not smaller than $0.95$ among the subsequent ($k$-wise) identified modes. The stabilisation diagrams for NExT-LF, NExT-ERA, and SSI-CVA are reported in Figure˜4.
@@
As shown in Figure˜5, the results of NExT-LF and SSI-CVA exhibit markedly stronger stabilisation than those of the NExT-ERA case, with the difference being most evident in damping estimates $\zeta_{n}$. In particular, NExT-ERA yields virtually no modes that stabilise in damping. This behaviour is consistent with the identified values of $\omega_{n}$ and $\zeta_{n}$ reported in Tables˜1 and 2, respectively.
@@
The $\omega_{n}$ identified using NExT-LF, NExT-ERA, and SSI-CVA are in close agreement with the benchmark results reported in Dessena et al. (2022), with the largest discrepancy being $3.42\%$ for $\omega_{1}$ obtained via NExT-ERA. In contrast, the $\zeta_{n}$ in Table˜2 show noticeably larger differences. Specifically, the absolute error in $\zeta_{1}$ exceeds $50\%$ for both NExT-LF and NExT-ERA, with NExT-LF providing a smaller deviation than NExT-ERA, while SSI-CVA limits the absolute error to approximately $17\%$. For $\zeta_{2}$ and $\zeta_{3}$, the NExT-LF estimates deviate by about $7.3\%$ in absolute terms, while NExT-ERA produces larger errors, namely $11.4\%$ and $7.46\%$, respectively. Overall, SSI-CVA yields damping estimates that are most consistent with the reference identification, although the absolute error for $\zeta_{1}$ remains above $17\%$.
After discussing the identifications of $\omega_{n}$ and $\zeta_{n}$, the results for $\bm{\phi}_{n}$ are now considered. Consistently with the frequency estimates, the mode shapes identified from the output-only measurements show excellent agreement with the benchmark, with the diagonal MAC values being equal, or very close, to unity. Figure˜6 presents the $\bm{\phi}_{n}$ obtained via NExT-LF, NExT-ERA, and SSI from output-only data, overlaid with the reference $\bm{\phi}_{n}$ reported in Dessena et al. (2022). The identified $\bm{\phi}_{n}$ closely reproduce the benchmark results.
Figure 6: $\bm{\phi}_{n}$ of the XB-2 wing spar identified using the output-only methods, overlaid with the benchmark results (retrieved from Dessena et al. (2025a)).
@@ 4 Airbus Helicopters H135 Bearingless Main Rotor Blade
In order to generalise the findings of the previous sections, a second benchmark aeronautical structure should be considered. For this aim, an Airbus Helicopters H135 BMR blade is chosen. The Airbus Helicopters H135, shown in Figure˜7(a), is a lightweight rotorcraft capable of a maximum forward flight speed of 136 kts, a maximum range of 342 nm, and a maximum take-off mass of 3100 kg. In this work, the focus is on its BMR blade, which is made of the pitch control cuff with an elastomeric lead-lag damper, an aerofoil section, and a flexbeam. This arrangement allows for incorporation of conventional mechanical flapping, lagging, and torsional hinges of the blade into a single structural arrangement. Structurally, each blade component is made of composite materials: the flexbeam is manufactured from E-glass/913 unidirectional prepregs with a cruciform cross-section whose slits reduce torsional stiffness while providing fail-safe load paths via double-lug root attachments, whereas the pitch control cuff and aerofoil skin are made of carbon fibre reinforced polymer. The approximate total mass of a single BMR is $40$ kg, of which roughly $7.5$ kg are tuning masses for frequency adjustment and lead-lag bending moment reduction Bansemir and Emmerling (1999). More information on the Airbus Helicopters H135 BMR blade is available in Bansemir and Müller (1997); Kampa et al. (1999).
From a geometric perspective, the BMR blade has an equivalent chord of 0.288 m and spans 4.99 m from the hub attachment. The H135 BMR blade specimen considered here, see Figure˜7(b), is an unserviceable [^5] In the aircraft maintenance sense. item that was retired from service in relatively good condition with over 4000 hours of remaining service life. This specimen was dynamically tested via AVT within the former Cranfield University BladeSense project Weber et al. (2021). It is worth noting that the AVT considered here is not a test under the blade operating conditions, which will be in rotating conditions while attached to the rotor hub. Nevertheless, this setup can still be considered OMA, as the input data is unknown.
@@
Figure 7: (a) Airbus Helicopters H135 helicopter in Western Counties Air Operations Unit livery (UK – retrieved from Felce ) and (b) the test setup of the ambient vibration testing on the Airbus Helicopter H135 BMR blade (adapted from Weber et al. (2021)).
The AVT was carried out in the aeroelastic laboratory of the Sir Peter Gregson Aerospace Integration Research Centre at Cranfield University, under controlled temperature and humidity conditions. The stationary, i.e. not revolving, blade is clamped at its end and instrumented with 9 nine triaxial accelerometers (A1-9), recording in-plane ($z$-axis) and out-of-plane transverse ($y$-axis) acceleration response from ambient vibration666xx-axis accelerations are recorded from A9, but are disregarded in this study.. The types and model of accelerometer are shown in Figure˜7(b) and their layout and spanwise position in Figure˜8. The locations of the accelerometers along the span in Figure˜8 are determined graphically from the Figures in Weber et al. (2021). Additionally, note that   highlights the 4 PCB Model 356B18 accelerometers used, while   and   identify, respectively, the 3 PCB Model 356A16 and 2 PCB Model 356A45 accelerometers. The use of three different accelerometer models was driven by laboratory availability rather than by a specific technical requirement; all three types provide adequate sensitivity for the low-amplitude ambient vibrations measured on this BMR blade across the frequency range of interest.
@@
18 acceleration time series are recorded from the 9 accelerometers at a sampling frequency $f_{s}=$ 2560 Hz for 600 s, using a similar setup to what is described for the XB-2 spar in Section˜3, involving a National Instruments cDAQ-9178 and then an in-house LabVIEW programme. Prior to identification, the signals are decimated to 256 Hz, trimmed, detrended, and bandpass-filtered between 0.4 and 100 Hz to exclude potential low- and high-frequency drifts with the bandpass [^7] https://uk.mathworks.com/help/signal/ref/bandpass.html built-in MATLAB function. From this, a PSD of the 18 channels and the ANPSD are found to qualitatively assess the quality of the response. The PSDs and ANPSD, shown in Figure˜9 are calculated using Welch’s method over 2048 discrete points with a Hamming window of length 2048 with an overlap of 256.
Figure 9: PSDs and ANPSD from the signals recorded from accelerometers A1-9 on the Airbus Helicopter H135 BMR blade.
As is clear from Figure˜9, this case poses a tougher identification challenge than the XB-2 spar, as the AVT amplitudes are much smaller than those produced by shaker excitation, resulting in lower modal amplitudes. As for the other cases, the identification procedure starts by obtaining the IRF from NExT, discretised over 4098 points, using two reference channels, A1 in $y$ and A4 in $z$ directions. This is necessary in this case, as using a reference channel for each direction increases the likelihood of obtaining accurate estimates of the modal parameters Caicedo (2011). The obtained IRFs are fed directly to ERA, while they are transformed to the frequency domain via an FFT discretised over 2048 points for use with the LF. The SSI-based identification is carried out directly from the time-domain data.
The identification is carried out in a given order range $k\in[22,130]$ and stabilisation diagrams are used to exclude spurious and non physical modes. For this aim, stabilisation parameters are defined. For a fair comparison, these remain unchanged across NExT-LF, NExT-ERA, and SSI-CVA. The stabilisation assessment is carried out on $f_{n}\in[0,100]\,$ Hz and $\zeta_{n}\in[0.003,0.03]$. Frequency stability is enforced with a relative tolerance of $0.5\%$ (i.e., $\Delta\omega_{n}/\omega_{n}\leq 0.005$), while damping stability is checked with a $5\%$ tolerance (i.e., $\Delta\zeta/\zeta\leq 0.05$). An additional absolute criterion on frequency variation is also applied, $|\Delta\omega_{n}|\leq\epsilon$, with $\epsilon=0.1$. A mode is accepted as stable only if these criteria are satisfied for three consecutive identifications. The consistency of $\bm{\phi}_{n}$ is enforced by requiring a MAC not smaller than $0.95$ among the subsequent stable identified modes. The stabilisation diagrams obtained from the NExT-LF, NExT-ERA, and SSI-CVA identification reported above are shown in Figure˜10.
@@
The extracted modal parameters, $\omega_{n}$ and $\zeta_{n}$, are reported in Table˜3 with results from the literature. These include results from AOMA Mugnaini et al. (2022), containing also $\bm{\phi}_{n}$, and Weber et al. (2021), including $\omega_{n}$ from a calibrated FEM and $\omega_{n}$ and $\zeta_{n}$ from a separate experimental configuration, based on a series of GVTs conducted using a random-on-random excitation waveform. Please note that in Table˜3, flapping refers to out-of-plane bending (within the xxyy plane), lagging to in-plane bending (within the xxyy plane), pitching to torsional rotations about the $x$-axis, and coupled to a mode exhibiting a combination of two or more of the dominant motions described above.
@@
It should be noted that the relative differences reported in Table˜3 are computed with respect to the AOMA results in Mugnaini et al. (2022) , as these represent the most directly comparable OMA-to-OMA reference for the same specimen and test configuration, and constitute the only available source from which reference mode shapes can be retrieved. The experimental values reported in Weber et al. (2021) originate from a different test setup (GVT with random-on-random excitation) and are therefore retained as an external validation benchmark rather than as the primary comparison baseline. It should also be noted that a change in testing conditions (e.g. boundary conditions, excitation type) can influence the identified modal response, which further motivates the use of results obtained under the same test configuration as the primary reference. From Table˜3, it is important to note that NExT-LF identifies two additional modes within the frequency range of interest, compared to the previous identification reported for AOMA Mugnaini et al. (2022) and the experimental results in Weber et al. (2021). These modes are located at $19.69~\mathrm{Hz}$ and $59.14~\mathrm{Hz}$, respectively. In addition, SSI-CVA identifies a further mode at $68.09~\mathrm{Hz}$. As shown later, these modes are all coupled in the $z$- and $y$-axis displacements, with the suspicion that the 2nd coupled mode may also include a torsional component; however, this cannot be fully confirmed due to the adopted instrumentation layout. Nevertheless, the coupled character of these modes provides a plausible explanation for why they are not identified by the simplified FEM reported in Weber et al. (2021). Furthermore, it is worth noting that NExT-LF finds more modes than NExT-ERA (13 vs 9), and the same number as SSI-CVA.
@@
In terms of the identified natural frequencies $\omega_{n}$, the agreement between the different output-only identification methods is generally very good. With respect to the reference values obtained via AOMA Mugnaini et al. (2022), the relative differences remain below approximately $3\%$ for all coincident modes, the largest deviation being observed for NExT-ERA at the 10th mode. A similar level of agreement is observed when compared with the experimental results reported in Weber et al. (2021). The consistency of the identified $\omega_{n}$ across NExT-LF, NExT-ERA, and SSI-CVA confirms the robustness of the proposed framework in terms of frequency estimation, even for higher-order flapping and lagging modes.
@@
However, the same level of consistency is not observed for damping ratios $\zeta_{n}$. In contrast to $\omega_{n}$, the identified $\zeta_{n}$ values exhibit a significantly larger dispersion between methods. In particular, relative differences with respect to AOMA Mugnaini et al. (2022) frequently exceed several tens of a percentage and, in some cases, are considerably larger, especially for lower-frequency modes. This behaviour is expected, as damping estimation in operational modal analysis is well known to be more sensitive to noise, limited measurement duration, and modelling assumptions than frequency estimation Phani and Woodhouse (2007). Moreover, small absolute differences in $\zeta_{n}$ can translate into large relative deviations when the reference damping level is low. This amplification effect is well documented in the OMA literature Phani and Woodhouse (2007) and has been recently confirmed for aerospace-grade structures in Zhu et al. (2025), where the sensitivity of output-only damping estimates to record length, noise floor, and spectral leakage is quantified.
@@
Overall, the results indicate that while the identification of $\omega_{n}$ is highly reliable and consistent among all the approaches considered, the estimation of $\zeta_{n}$ remains more method-dependent. However, to fully assess the performance of NExT-LF, the identified $\bm{\phi}_{n}$ must also be examined. To this end, Table˜4 reports the diagonal MAC values between the $\bm{\phi}_{n}$ identified via NExT-LF and those obtained in Mugnaini et al. (2022).
@@
The MAC results in Table˜4 further contextualise the comparisons in identified $\omega_{n}$ and $\zeta_{n}$. In particular, NExT-LF achieves a very high correlation for several dominant modes (e.g., Modes 1, 3, 4, 8, and 9, with MAC values close to unity), supporting the strong agreement observed for $\omega_{n}$ and indicating that the associated $\bm{\phi}_{n}$ are consistently retrieved. At the same time, the MAC values highlight that the increased modal content identified by NExT-LF comes with a non-uniform level of shape fidelity across all modes: some modes exhibit reduced correlation with respect to Mugnaini et al. (2022), notably Mode 5 (MAC =0.77=0.77) and, more markedly, Mode 7 (MAC =0.44=0.44), as well as very low agreement for Mode 10 (MAC =0.16=0.16) and Mode 14 (MAC =0.09=0.09). This suggests that, although NExT-LF is effective in extracting a larger set of modes than NExT-ERA, a subset of the identified shapes may be more sensitive to the limited instrumentation and to the specific characteristics of the identification procedure, leading to a weaker correspondence with the AOMA reference in Mugnaini et al. (2022). Conversely, SSI-CVA shows consistently high MAC values for the modes that can be compared (e.g., Mode 7 and Mode 10), indicating stronger agreement in those cases, albeit without the same increase in the number of extracted modes relative to NExT-ERA. For completeness, the 14 identified $\bm{\phi}_{n}$ are shown in Figure˜11. Note that in the cases where the MAC is more than 0.95, the NExT-LF-identified $\bm{\phi}_{n}$ are shown; otherwise, $\bm{\phi}_{n}$ from Mugnaini et al. (2022) are preferred, with an exception made for mode 12 (3rd coupled), which is retrieved from the SSI-CVA identification.
@@
All $\bm{\phi}_{n}$ shown in Figure˜11 are consistent with the dominant motions reported in Table˜3. Nevertheless, the trajectories associated with the two pitching modes (Figures˜11(g) and 11(n)) cannot be directly inferred from the corresponding graphs, as rotational degrees of freedom are not measured on the blade. Therefore, these modes are classified as pitching based on the evidence suggested by the FE analyses reported in Weber et al. (2021). In addition, the three newly identified modes in the present campaign – namely, the coupled modes (Figures˜11(f), 11(k) and 11(l)) – do not exhibit a clear dominant displacement direction. For this reason, they are labelled as coupled, at least in terms of combined lagging and flapping components.